OpenStudy (anonymous):

What is S10 for 1+2+4+8+...?

OpenStudy (aravindg):

you mean sum of 10 terms?

OpenStudy (anonymous):

If these answers choices relate then yes: 256.5 511 1,023 2047

OpenStudy (saifoo.khan):

$10

OpenStudy (saifoo.khan):

| was missing. @AravindG

OpenStudy (aravindg):

Well this is a geometric progression

OpenStudy (aravindg):

Can you find the common ratio ?

OpenStudy (anonymous):

*2?

OpenStudy (aravindg):

yes :)

OpenStudy (aravindg):

Now \[S_n=\dfrac{a(r^n-1)}{(r-1)}\]

OpenStudy (aravindg):

This is the sum formula .Put r=2 a=1

OpenStudy (aravindg):

n=10 .You will get sum

OpenStudy (anonymous):

using the geometric formula, find out the 10th term:\[a_n=a(2)^{n-1}\]\[a_{10}=1(2)^{10-1}\]\[a_10=2^9\]\[a_10=512\]now using the formula for the sum of a finite geometric series, find the sum:\[{a_1(1-r^n)}\over1-r\]\[{1(1-2^10)}\over1-2\]\[(1-1024)\over-1\]\[-1023\over-1\]\[S_10=1023\]

OpenStudy (aravindg):

@yummydum you didnt leave anything to do for the asker ? :/

OpenStudy (anonymous):

wow im bad at this..

OpenStudy (saifoo.khan):

@AravindG : We have so many bad users here.. :l

OpenStudy (aravindg):

Thats sad :/

OpenStudy (anonymous):

han! aur aik ho tum, kaminay kaheen k! :P

OpenStudy (anonymous):

i leave..

OpenStudy (anonymous):

0_0...

OpenStudy (aravindg):

@xkylex Dont mind our conversation..Did you understand how to do it?

OpenStudy (anonymous):

I understood perfectly. Thank You all!

OpenStudy (aravindg):

yw :)